OpenStudy (anonymous):

Find the total interest paid on the given amortized loan. Arianne bought new furniture for $4900. she paid a $850 down payment and financed the rest for 5 years at an interest rate of 8.3%.

OpenStudy (anonymous):

I dont know. But. You are hot.

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

lol. thanks..

OpenStudy (anonymous):

Are you asian?

jimthompson5910 (jim_thompson5910):

"Arianne bought new furniture for $4900. she paid a $850 down payment" how much is financed? ie, how much money is she loaned?

OpenStudy (anonymous):

4050 is the loan

OpenStudy (anonymous):

xilly why you ignore me?

OpenStudy (anonymous):

you broke my heart </3

OpenStudy (anonymous):

no im not asian.

OpenStudy (anonymous):

want to be my girlfriend? ;)

OpenStudy (anonymous):

i have a boyfriend already c;

jimthompson5910 (jim_thompson5910):

correct, 4900-850 = 4050 is the amount loaned out now you'll use this formula P = L[c(1 + c)^n]/[(1 + c)^n - 1] where L = 4050, c = 0.083/12 = 0.00691667, and n = 12*5 = 60 or you can use this calculator http://www.bankrate.com/calculators/mortgages/mortgage-calculator.aspx it says "mortgage", but it can be used for any amortized loan

jimthompson5910 (jim_thompson5910):

absolutemaster this isn't the place for this sort of thing

OpenStudy (anonymous):

82.70?

jimthompson5910 (jim_thompson5910):

I'm getting $82.70 as well

jimthompson5910 (jim_thompson5910):

that's her monthly payment

jimthompson5910 (jim_thompson5910):

how much will she pay back in total?

OpenStudy (anonymous):

Then what do i do ?

jimthompson5910 (jim_thompson5910):

hint: 5 years = 60 months since 5*12 = 60

OpenStudy (anonymous):

so multiply the answer by 60?

jimthompson5910 (jim_thompson5910):

correct, to find the total paid back

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

4962.13

jimthompson5910 (jim_thompson5910):

you did 60*82.70 right?

OpenStudy (anonymous):

well i just hit the answer key*60

OpenStudy (anonymous):

82.70*60=4962

jimthompson5910 (jim_thompson5910):

yeah it should be 60*82.70 = 4962

jimthompson5910 (jim_thompson5910):

so how much interest does she pay?

OpenStudy (anonymous):

62?

OpenStudy (anonymous):

A)930.14 B)921.33 C)927.13 D)912.13 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hmm I'm getting 4962 - 4050 = 912 idk how they're getting 912.13

jimthompson5910 (jim_thompson5910):

there seems to be a rounding error somewhere

jimthompson5910 (jim_thompson5910):

oh I see now

OpenStudy (anonymous):

oh.. i was subracting by 4900

jimthompson5910 (jim_thompson5910):

the monthly payment is really 82.702109022983 if you don't round

jimthompson5910 (jim_thompson5910):

82.702109022983*60 = 4,962.126541379

jimthompson5910 (jim_thompson5910):

4,962.126541379 - 4050 = 912.126541379001 which rounds to 912.13

OpenStudy (anonymous):

ok i get it , thank you!!

jimthompson5910 (jim_thompson5910):

np